OpenStudy (anonymous):

A ball traveling with an initial momentum of 2.5 kg·m/s bounces off a wall and comes back in the opposite direction with a momentum of -2.5kg·m/s. a. What is the change in momentum of the ball? b. What impulse would be required to produce this change?

OpenStudy (anonymous):

Change is (2.5) -(-2.5) = 5.0 F = m (dv/dt) F dt = impulse = m (dv) = d(mv) = 5 kg-m/s